OpenStudy (bamboo888):

can you help with 6 questions about writing algebraic expressions? Thanks!

OpenStudy (bamboo888):

Directions: Define a variable. Then write each phrase as an algebraic expression 1. five times the number of books in the library 2. two more than seven times Lynn's age 3. The distance around a basketball, or circumference, is about three times the circumference of a softball. Define a variable and write an expression to represent the circumference of a basketball. 4. A camp leader figures that she needs one tent for every three campers, plus a tent for herself. Define a variable and write an expression to represent the number of tents needed. Table needed for Q5&6: Country: Medals: Country: Medals: Country: Medals: USA m Great Britain 65 France 34 China 88 Australia a South Korea 28 Russia 82 Germany 44 Italy i 5. Write an expression to show how many fever medals Australia won than Great Britain. 6. The number of medals won by the US is 20 more than three times the number of medals won by Italy. Write an algebraic expression to represent this statement.

OpenStudy (bamboo888):

Will medal and fan!

OpenStudy (yoloshroom):

Hi there, for Number 1 we can say "X = The Number Of Books In The Library" then we can asnwer number 1 as, 5*x or 5x or 5(x) - all mean 5 times X

OpenStudy (anonymous):

1. y=5n books 2. y=7n+2 years old 3.

OpenStudy (yoloshroom):

2+ 7x for the second one we need to define our variable again, so we can say "Lynns Age is equal to X" so the answer would be 2 + 7x which translates to two plus (more than) 7 times X

OpenStudy (yoloshroom):

3) We can set "Circumference Of Basketball" equivalent to B, and the"Circumference Softball" equivalent to S. so, B=3S

OpenStudy (yoloshroom):

@Brill Brain just died, finisht htis off for me?

OpenStudy (anonymous):

K sure hold on

OpenStudy (anonymous):

4) If x is the number of campers this can be defined by \[\frac{ 1 }{ 3x }+1=y\]

OpenStudy (anonymous):

5) if australia's number of medals is a then the equation is 65-a=y

OpenStudy (bamboo888):

What about the last one? I was soo stuck on that one. And, can you explain a little more on Q4? @Brill

OpenStudy (anonymous):

for 6) The number of medals won by Italy is i, if the US has 20 more than 3 times i it is, y=20+3i

OpenStudy (anonymous):

And sorry about number 4, I typed that in wrong, it is \[\frac{ 1 }{ 3 }x+1=y\]

OpenStudy (anonymous):

It is that way because for every 3 campers there is 1 tent